The product of two consecutive positive odd integers is one less than three times their sum. Find the integers.
can someone show me how to set this up im so confused

Found 2 solutions by scott8148, checkley77:
Answer by scott8148(6628)

consecutive odd integers are two units apart (3 and 5 , 11 and 13 , etc.)

n(n+2) + 1 = 3(n + n + 2)

n^2 + 2n + 1 = 6n + 6

n^2 - 4n - 5 = 0

(n - 5)(n + 1) = 0

n = 5 , n+2 = 7

Answer by checkley77(12844)

X & X+2 ARE THE 2 CONSECUTIVE INTEGERS.
X(X+2)=3(X+X+2)-1
X^2+2X=3(2X+2)-1
X^2+2X=6X+6-1
X^2+2X-6X-5=0
X^2-4X-5=0
(X-5)(X+1)=0
X-5=0
X=5 ANS.
5+2=7 ANS.
X+1=0
X=-1 ANS.
-1+2=1 ANS.
PROOF:
5*7=3(5+7)-1
35=3*12-1
35=36-1
35=35